# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

### 22 Mar 2026 v2

@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$ (see Section LABEL:ap:eq:marlikcond of the supplementary materials). The differential-shrinkage $g$ prior can be obtained by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. Indeed, note that the standard $g$ prior is then obtained by further setting $g_{1}=g_{2}=\cdots=g_{p_{\boldsymbol{\gamma}}}=g$.
@@ 4.1 Tail behavior
This implies that the our prior has heavy (polynomial) tails and point estimators derived from our procedures is robust, in the sense of having bounded influence in the case of likelihood-prior conflict. However, note that the contour plots associated with the DP mixture of block $g$ priors are not elliptical, unlike those of the standard $g$ prior (please see Section LABEL:ap:scatterplotsbivariagte of the supplementary materials). These non-elliptical contours enable non-uniform shrinkage by the DP block $g$ priors, which in turn is what enables them to skirt the conditional Lindley paradox.
@@ 4.2 Information consistency of Bayes factors
The proof is included in Section LABEL:ap:informationconsistency of the supplementary materials. A slightly simpler condition that applies to DP mixtures of block $g$ priors is the following:
@@
See Section LABEL:ap:informationconsistency of the supplementary materials. This is the same condition in Theorem 2 of Liang et al. (2008). Hence, this result just indicates that any mixing distribution for $g$ that leads to information-consistent Bayes factors under a standard mixture of $g$ priors also leads to information consistent Bayes factors under the DP mixture of block $g$ priors.
@@ 4.3 Information consistency of block structures
The proof can be seen in Section LABEL:ap:cluseringprop of the supplementary materials.
@@ 4.4 Conditional Lindley paradox
The proof, which can be found in Section LABEL:ap:condLindleyparadox of the supplementary materials, only requires that the prior on the partition $\rho$ puts positive probability on at least one refinement of the true partition $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$. This is trivially true for the Dirichlet process, whether its concentration parameter is fixed or treated as unknown and given a prior distribution. In fact, this is a very mild requirement that is also satisfied by the Bayes factor constructed under the global-local $g$ prior, and one that does not require that the partition model be able to identify the true partition $\rho_{0}$. On the other hand, orthogonality plays a key role in the proof of this theorem, just like block orthogonality is key to similar proofs in Som (2014). Simulation studies suggest that the result hold in the non-orthogonal case.
@@ 4.5 Model selection consistency
Model selection consistency refers to the ability of the procedure to choose the correct model as the sample size grows. DP mixtures of block $g$ priors are model selection consistent in the fixed $p$ regime (see Section LABEL:ap:modelselconsistency of the supplementary materials).

### 12 May 2026 v3

@@ 3 A new class of priors: Dirichlet process mixtures of block $g$ priors
where $\boldsymbol{\Omega}_{\boldsymbol{\gamma}}=\boldsymbol{I}_{n}+\boldsymbol{X}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\boldsymbol{G}_{\boldsymbol{\gamma}}^{1/2}\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}$ (see Section A of the supplementary materials). The differential-shrinkage $g$ prior can be obtained by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left\{\boldsymbol{X}_{\boldsymbol{\gamma}}^{T}\boldsymbol{X}_{\boldsymbol{\gamma}}\right\}^{-1}$. Indeed, note that the standard $g$ prior is then obtained by further setting $g_{1}=g_{2}=\cdots=g_{p_{\boldsymbol{\gamma}}}=g$.
@@ 4.1 Tail behavior
This implies that the our prior has heavy (polynomial) tails and point estimators derived from our procedures is robust, in the sense of having bounded influence in the case of likelihood-prior conflict. However, note that the contour plots associated with the DP mixture of block $g$ priors are not elliptical, unlike those of the standard $g$ prior (please see Section B of the supplementary materials). These non-elliptical contours enable non-uniform shrinkage by the DP block $g$ priors, which in turn is what enables them to skirt the conditional Lindley paradox.
@@ 4.2 Information consistency of Bayes factors
The proof is included in Section C of the supplementary materials. A slightly simpler condition that applies to DP mixtures of block $g$ priors is the following:
@@
See Section C of the supplementary materials. This is the same condition in Theorem 2 of Liang et al. (2008). Hence, this result just indicates that any mixing distribution for $g$ that leads to information-consistent Bayes factors under a standard mixture of $g$ priors also leads to information consistent Bayes factors under the DP mixture of block $g$ priors.
@@ 4.3 Information consistency of block structures
The proof can be seen in Section D of the supplementary materials.
@@ 4.4 Conditional Lindley paradox
The proof, which can be found in Section E of the supplementary materials, only requires that the prior on the partition $\rho$ puts positive probability on at least one refinement of the true partition $\rho_{0}=\{\mathcal{I}_{1},\mathcal{I}_{2}\}$. This is trivially true for the Dirichlet process, whether its concentration parameter is fixed or treated as unknown and given a prior distribution. In fact, this is a very mild requirement that is also satisfied by the Bayes factor constructed under the global-local $g$ prior, and one that does not require that the partition model be able to identify the true partition $\rho_{0}$. On the other hand, orthogonality plays a key role in the proof of this theorem, just like block orthogonality is key to similar proofs in Som (2014). Simulation studies suggest that the result hold in the non-orthogonal case.
@@ 4.5 Model selection consistency
Model selection consistency refers to the ability of the procedure to choose the correct model as the sample size grows. DP mixtures of block $g$ priors are model selection consistent in the fixed $p$ regime (see Section F of the supplementary materials).